Mondi Company produces party boxes that are sold in bundles of 1000 boxes. The market is highly competitive with boxes currently selling for R100 per thousand. The company has a total and marginal cost curve given by: TC= 3 000 000+0.001Q² MC= 0.002Q Q is measured in thousand boxes bundles per year a) Determine Mondi's profit maximizing quantity? b) Calculate if the firm is earning a profit or loss? c) Based on the analysis above, should Mondi Company operate or shut down in the shortrun?
1
Expert's answer
2021-10-25T09:16:55-0400

(a)

"TC=3,000,000+0.001Q^2"

"MC=0.002Q"

"TR=P\times Q"

"=100Q"

"MR=100"

At profit maximizing quantity:

"MC=MR"

"0.002Q=100"

"Q=50,000."

(b)

"TR=100\times50,000=5,000,000"

"TC=3,000,000+0.001(50,000^2)."

"=25,003,000,000"

Profit/Loss "=TR -TC"

"=5,000,000 - 25,003,000,000"

"=-20,003,000,000."

Thus, the firm is earning a loss.

(c)

In the short run, the firm should shut down because from the analysis, it is not earning revenue to cover its variable costs.